You want to estimate the mean weight of quarters in circulation. A sample of 60 quarters has a mean weight of 5.625 grams and a standard deviation of 0.063 gram. Use a single value to estimate the mean weight of all quarters.​ Also, find the​ 95% confidence interval for the average weight of all quarters.

The estimate for the mean weight of all quarters is ___ grams?


1
Expert's answer
2021-06-01T17:27:56-0400

The estimate for the mean weight of all quarters is "5.625" grams.


The critical value for "\alpha=0.05" and "df=n-1=60-1=59" degrees of freedom is "t_c=z_{1-\alpha/2, n-1}=2.000995." The corresponding confidence interval is computed as shown below:


"CI=(\bar{x}-t_c\times\dfrac{s}{\sqrt{n}}, \bar{x}+t_c\times\dfrac{s}{\sqrt{n}})""=(5.625-2.000995\times\dfrac{0.063}{\sqrt{60}},"

"5.625+2.000995\times\dfrac{0.063}{\sqrt{60}})"

"=(5.608725, 5.641275"

Therefore, based on the data provided, the 95% confidence interval for the population mean is "5.609<\mu<5.641," which indicates that we are 95% confident that the true population mean "\mu" is contained by the interval "(5.609, 5.641)."